OpenStudy (anonymous):

4.A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (aaronq):

|x − 0.525| = 0.01 So we have 2 scenarios, where the two furthest points lie \(\pm\)0.01 from 0.525 x − 0.525 = 0.01 -(x − 0.525) = 0.01 x = 0.01+0.525 -x + 0.525 = 0.01 x =0.535 -x = 0.01-0.525 -x = -0.515 x = 0.515